OpenStudy (anonymous):

help me !!! Given the function g(x) = −3x + 4, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is smaller than the value of g(4). The value of g(−2) is the same as the value of g(4). The values of g(−2) and g(4) cannot be compared. The value of g(−2) is larger than the value of g(4).

OpenStudy (anonymous):

it's the last one coz -2 into - it will +ve so it greater than 4

OpenStudy (campbell_st):

just substitute x = -2 then x = 4 into the equation and evaluate g(-2) = -3 x (-2) + 4.....= g(4) = -3 x (4) + 4 ... =

OpenStudy (anonymous):

thank you moha_10

OpenStudy (anonymous):

ur welcome @ivandelgado